a tennis ball is hit vertically upwards with a velocity of 20m/s.find the time taken for the ball to reach its highest point


1
Expert's answer
2021-12-23T10:49:04-0500
"v=v_0+gt,""0=v_0+gt,""t=-\dfrac{v_0}{g}=-\dfrac{20\ \dfrac{m}{s}}{-9.8\ \dfrac{m}{s^2}}=2.04\ s."